You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100614 (15 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June 13-14

Observer: V. Testa, K.Boutsia
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O. Kuhn
Instrument: LUCIFER1 - LBC

Summary

Clear sky - seeing <1.0''.
Problems with the focus of LBC probably due to a misconfiguration of the dofpia

Data Summary

Project Objects Filter ExpTime Notes
Mannucci Q1422 210zJHK H 1hr15min arc- flat- tell
Mannucci Q1422 imaging Ks 1hr  
Mannucci Q1422 imaging J 1hr  

Details

02:20 - Opening the dome - starting interface

02:30 - test Flats to monitor sky brightness 0001-0021

02:35 - Sky Flat fields

in the Ks band - imaging
  • 2secs / 11000-9000counts 0022-0031
  • 2secs / 7000-5000counts 0033-0042
  • 2secs / <4000counts 0043-0052

in the H band
  • 2secs / 12000-10000counts 0054-0063
  • 2secs / 4000-3000counts 0074-0083

in the J band - imaging
  • saturated 0032,0053
  • 2secs / 10000-8000counts 0064-0073
  • 2secs / <3000counts 0084-0093

03:00 - Pointing Star near the field (WT10_303) 0094-0095
  • No need to do high elevation - low elevation any more because it has been stable

03:18 - Persson standard BS9153 for collimation (POSITION,ACTIVE)

03:24 - Collimation - preset to Mannucci Q1422 acquisition script for imaging
  • Q1422_imaging_acq.txt (script used)

03:27 - Mannucci Q1422 science script for imaging - Ks
  • Q1422_Ks_sci.txt 0096-0135
  • 15DIT x 6NDIT x 40 exp (with 20.0'' jitter)

04:50 - Mannucci Q1422 science script for imaging - J
  • first we repeat acq script to center field after jittering
  • Q1422_J.txt 0136-0175
  • 30DIT x 3NDIT x 40 exp (with 20.0'' jitter)

06:10 - Mannucci Q1422 - MOS acquisition 00176-0181

06:40 - Mannucci Q1422 SCI script H_4_ext 0182-0197
  • N1.8 - 210_zJHK- K / 300secx1x16offset (seeing <1.0)
  • During the 9th iteration 0190 gives flexure compensation error
  • integration is not interrupted
  • we switch OFF and ON flexure compensation (without committing telescope)
  • in the middle of 10th iteration 0191 flexure compensation is working again

08:08 - Calibration Unit ON - ARC and FLAT scripts with Mannucci
  • ARCS Ar OFF 0198
  • ARCS Ar ON 0199
  • ARCS Xe OFF 0200
  • ARCS Xe ON 0201
  • FLAT Halo1 OFF 10sec 0202-0204
  • FLAT Halo1 ON 10sec 0205-0207
  • We have doubled the exposure time for the flat aquisition

08:20 - Pointing correction (ACT0262) 0208-0210

08:29 - Telluric aquisition in the H band after correcting pointing 0211

08:37 - Telluric SCI script (mannucci_tell_int_Q1422_H.txt) 0212-0219

08:42 - Telluric aquisition in the H band again because telluric was offset 0220-0222

08:46 - Telluric SCI script (mannucci_tell_int_Q1422_H.txt) 0223-0230
  • we have changed integration to 10secDITx2NDIT

08:52 - SWITCH TO LBC

09:17 - Slew to WT_10_302

09:24 - Slew to NGC 6791, crazy pupils, moved 20' S and redo dofpia

09:43 - Red immediately out of focus

09:52 - stopped and redo dofpia

10:08 - Using a wrong (old) config file for dofpia finished the OB but almost useless

10:50 -Using the new config file things went much better repeated NGC6791_WD_LONG_1

Sky already bright, did not finish the OB

11:20 - Move to Flat Field (modified OB to slew to the FF spot because it won't dither anymore if dec=-90 and not presetting) Must do this for every FF OB and for every run.....how nice!

11:25 - SkyFlat_VV 5s

11:40 - SkyFlat Bi 0.5s

11:45 - End of night....pheewwww!!!!!

Send Dark sequence on Lucifer after parking the telescope

-- VincenzoTesta - 14 Jun 2010
Topic revision: r6 - 15 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback